### v2

@@ 3.1.1 Statistical Models
Prior to machine learning and deep learning, traditional statistical models, which laid the foundation for analyzing sequential data, were commonly utilized for time series forecasting. Exponential smoothing was introduced by [15], [73], and [189] as a method to forecast future values using a weighted average of past data. This method operates by computing an exponential weight decay of historical data, assigning greater weight to more recent data. The Autoregressive Integrated Moving Average (ARIMA) model was formalized through the work of George Box and Gwilym Jenkins in their book “Time Series Analysis: Forecasting and Control [11].” The ARIMA model predicts future values by leveraging the autocorrelation in data and is composed of three main components. AutoRegressive (AR) uses a linear combination of past values to predict the current value. Moving Average (MA) employs a linear combination of past error terms to predict the current value. Integrated (I) removes non-stationarity (the property where mean and variance change over time) by differencing the data to achieve stationarity. While exponential smoothing models are advantageous for real-time data analysis due to their simplicity, ARIMA models are better suited for capturing complex patterns, making them ideal for long-term forecasting. The SARIMA model extends ARIMA by incorporating seasonal differencing, along with seasonal autoregressive and moving average components, allowing it to effectively model and predict data with regular cyclical patterns. These statistical models are based on specific assumptions and are simple, intuitive, and useful for identifying basic patterns in data.
@@ 3.1.2 Machine Learning Models
3.2 Traditional Deep Learning Models
@@ 3.2.1 MLPs: The Emergence and Constraints of Early Artificial Neural Networks
Recurrent Neural Networks (RNNs) [74] emerged, opening up new possibilities for processing time series data. RNNs are specialized models designed to process sequential data, such as time series data, and were utilized to overcome the limitations of MLPs. The structure of an RNN consists of an input layer, a hidden layer, and an output layer, and it uses the output from previous time steps(hidden layer output) as the input for the current time step. This structural feature allows RNNs to naturally model the dependencies in data over time. The hidden state preserves past information by storing data from the previous time step and passing it on to the next time step, thereby maintaining historical context. Additionally, it is parametrically efficient because the same weights are used for calculations across all time steps, regardless of the sequence length.
@@
To address the aforementioned issues with vanilla RNNs, researchers began developing various models. Notably, the Long Short-Term Memory (LSTM) [72] model was developed to address the long-term dependency issues of vanilla RNNs. With the cell state and gate mechanisms (input, output, and forget gates), LSTMs preserves crucial information over extended periods and the discards unnecessary information. The Gated Recurrent Unit (GRU) [30] emerged as a simpler alternative to LSTM, offering similar performance while using a more straightforward structure(update and reset gates) to manage information. These two models effectively addressed the vanishing gradient problem and sparked a boom in RNN-based models for time series analysis. Many subsequent RNN-based models used these two as their backbone, making them powerful tools for handling time series data until the advent of Transformers.
@@
The Dilated RNN [21] proposed Dilated Recurrent Skip Connections as a solution to various issues inherent in traditional vanilla RNNs. This approach enhanced parameter efficiency in long sequences and enabled the network to effectively learn complex dependencies across diverse temporal scales. The DA-RNN [143] attempted to effectively address long-term dependencies, which were challenging for existing methods, by employing a dual-stage attention mechanism. Utilizing an encoder with Input Attention and a decoder with Temporal Attention, it adaptively selected salient information from input features and temporal steps, thereby enhancing the accuracy and performance of time series predictions. Additionally, the MQ-RNN [188] combined the Sequence-to-Sequence neural network framework with Quantile Regression to provide probabilistic forecasts, thereby reflecting uncertainties across various scenarios. This model simultaneously generated multiple forecast horizons at each time step and employed the Forking-Sequences training method to enhance the stability and performance of the RNN.
@@ 3.2.3 CNNs: Extracting key patterns in time series data beyond just images
CNN and RNN Ensemble Models
Building on the popularity of traditional time series forecasting models like RNNs and the emerging interest in CNN models, hybrid CNN-RNN models began to emerge, combining the strengths of both architectures. Models that combined CNNs and LSTMs were particularly advantageous for simultaneously learning local patterns and long-term dependencies in time series data. The structure involved using CNNs to extract complex features from the time series data and LSTMs to learn temporal dependencies.
DeepAR [154] combined LSTM and CNN to generate probabilistic forecasts of future distributions rather than single predicted values. Additionally, it improved generalization performance by simultaneously learning and predicting multiple time series within a single model. DCRNN [97] modeled traffic flow as a diffusion process on a directed graph, capturing both temporal and spatial dependencies. The Diffusion Convolution leveraged bidirectional random walks on the graph to capture spatial dependencies within the traffic network, while the Gated Recurrent Units (GRU) modeled temporal dependencies. The introduction of the Diffusion Convolutional Gated Recurrent Unit (DCGRU) enabled the effective processing of temporal information. the TPA-LSTM [161] combined CNN and LSTM with an added attention mechanism, introducing the Temporal Pattern Attention(TPA) methodology to learn important patterns in time series data. Local patterns were extracted using CNN, while the combination of LSTM and the attention mechanism captured significant temporal patterns.
@@ 3.2.4 GNNs: Structurally modeling relationships between variables
Until the development of Transformers, GNNs consistently garnered interest, showing consistent yet slow growth compared to RNNs and CNNs. The development of Graph Convolutional Networks (GCNs) [89] marked another turning point for GNNs. GCNs learns node features through convolution operations on graph structures, and are powerful tools for processing graph data. This advancement positioned GCNs as a versatile solution for addressing a wide range of graph-based challenges.
@@
After the era of traditional statistical techniques and machine learning methodologies, traditional deep learning approaches such as RNNs, CNNs, and GNNs were widely used for time-series analysis and prediction over a long period. However, the advent of Transformers brought a revolutionary change to the AI field. Upon their introduction, Transformers outperformed all existing methods across various domains and continue to dominate as the leading foundation model. In the time-series domain, a similar trend was observed. Beginning with the introduction of the Logsparse Transformer [96], numerous transformer-based models emerged, achieving state-of-the-art performance. As a result, traditional deep learning methods experienced a period akin to the Ice Age. The remarkable historical TSF models mentioned above are organized chronologically in Fig. 8.
@@ 3.3 The Prominence of Transformer-based Models
Transformer [178], introduced by the Google Brain team, is an innovative model designed to perform complex sequence-to-sequence tasks in natural language processing. This model features an encoder-decoder structure and utilizes an attention mechanism to capture the relationships between tokens in the input sequence. As a result, Transformers have gained significant attention for replacing traditional RNN-based models, offering parallel processing capabilities and effectively addressing long-term dependency issues. Transformers’ success with sequential data naturally led to their extension into time series applications.
@@ 3.3.1 Transformer Variants
In recent years, substantial research has focused on transformer-based time series analysis, particularly long-term time series forecasting (LTSF), which has led to the development of various models [209]. The original Transformer model has several limitations when applied to long-term time series forecasting (TSF). These limitations include quadratic time and memory complexity, as well as error accumulation caused by the auto-regressive decoder design. Specifically, the time and memory complexity of self-attention increases quadratically with the length of the input sequence. This high complexity can become a computational bottleneck, especially for time series forecasting tasks that rely on long-term historical information. To address these issues, Transformer variants primarily focus on reducing time and memory complexity.
@@
Figure 9: Comparison of Full and Sparse Attention Mechanisms in Transformer
LogTrans [96] leverages the efficiency of Log Sparse attention mechanisms to reduce time and memory complexity to $O(LlogL)$, making it well-suited for long sequence forecasting tasks. This model assumes recent data is more relevant than older data, utilizing positional sparsity to focus attention only on relevant data. Reformer [90] enhances efficiency through locality-sensitive hashing and reversible residual layers, reducing memory and computational requirements to $O(LlogL)$ while maintaining performance on long sequences. Locality-sensitive hashing (LSH) approximates attention by clustering similar items and performing attention partially on these clusters. Informer [219] improves computational and memory efficiency to $O(LlogL)$ by utilizing a ProbSparse self-attention mechanism and a generative-style decoder. It effectively handles long sequences and captures dependencies across various time scales. Informer addresses the limitations of Reformer, which performs well only on extremely long sequences, and LogTrans, which uses heuristic positional-based attention. ProbSparse self-attention features dynamic query selection based on query sparsity measurement. Autoformer [191] integrates time series decomposition into the internal blocks of the Transformer and utilizes an auto-correlation mechanism to aggregate similar sub-series. Autoformer improves upon previous methods by addressing point-wise dependency and aggregation issues. It employs a decomposition architecture to handle complex temporal patterns by separating seasonal and trend components. Pyraformer [110] introduces a novel pyramid attention mechanism designed to efficiently model long-term dependencies in time series data. By using a hierarchical structure to progressively reduce sequence length, Pyraformer significantly lowers computational complexity compared to traditional Transformer models. Fedformer [220] combines Fourier-enhanced decomposition blocks with frequency domain sparse representations, integrating seasonal-trend decomposition from Autoformer and offering a more effective representation of time series data. The sparse representation basis can be Fourier-based or Wavelet-based, encompassing a subset of frequency components, including both high and low frequencies. Additionally, attempts have been made to overcome the limitations of normalized Transformers, which often produce indistinguishable attention patterns and fail to capture significant temporal dependencies. Non-stationary Transformer [111] identifies this issue as excessive normalization and addresses it with De-stationary Attention, which allows for the use of non-normalized data during QKV computations.
@@ 3.3.2 Limitation of Transformer-based Models
Finite Context Window
The efficiency of self-attention is attributed to its ability to densely route information within a context window, enabling the modeling of complex data. However, unlike models such as RNNs or State Space Models (SSMs), Transformers have a fundamental limitation in modeling data beyond a finite window.
@@ 3.4 Uprising of Non-Transformer-based Models
Contemporary TSF tasks increasingly require the prediction of diverse multivariates and long sequences. Unlike earlier tasks that involved relatively short sequences, the limitations of transformer-based models have become more pronounced when dealing with long sequences. Researchers adhering to the Transformer’s philosophy have begun focusing intensively on two aspects to address these issues: reducing computational complexity and enhancing long-term dependency learning. Conversely, some researchers have reverted to non-transformer methodologies, renewing interest in traditional deep learning models such as RNNs, CNNs, GNNs, and MLPs. Non-transformer-based models have the potential to overcome the limitations of Transformers, and research has begun to leverage these strengths to move beyond the hegemony of the Transformer backbone.
RNN-based models sequentially process input data, retaining the previous state at each time step to predict the next state. This enables the natural modeling of long-term dependencies and helps overcome the limitations of Transformers by leveraging time-ordering learning. They require relatively less data and memory, making them less restrictive in thier application. Additionally, their ability to handle real-time streaming data enhances their versatility across various applications.
@@
This claim gained further traction through subsequent studies, leading to a more diversified interest in backbone models across various architectures. The emergence of this model sparked skepticism regarding the performance of Transformer-based models and served as a catalyst for the exploration of various alternative architectures. The era of Transformer prevalence has come to an end, ushering in a renaissance in time series research, characterized by a reevaluation of traditional DNNs, the emergence of novel models like Mamba, and new approaches such as the utilization of LMMs.
@@ 4.1 Overcoming Limitations of Transformer
As previously mentioned in 3.4, the performance of simple linear models in LTSF-Linear [209] surpassing traditional transformer-based models has raised doubts about the effectiveness of Transformers. However, in the field of NLP, Transformers still demonstrate superior performance in handling long-term dependencies in sequential data compared to other models [139]. This observation suggests that while Transformers have great potential, researchers have not fully leveraged their capabilities in time series analysis.
@@ 4.1.1 Patching Technique
Transformers were originally developed for natural language processing (NLP), and applying them to time-series analysis requires appropriate adjustments to fit the domain of time series. Different from the rich semantic information carried by individual word tokens in NLP, individual data points in time series are often similar to their neighboring values, lacking substantial information [134]. Therefore, the point-wise attention mechanism of traditional models fails to consider the broader context or patterns that may span multiple consecutive time steps and only calculates attention for individual time steps, which makes it difficult to capture the characteristics of time series data. For this reason, considering the surrounding context along with a single time point provides more information in time series data.

### v3

@@ 3.1.1 Statistical Models
Prior to machine learning and deep learning, traditional statistical models, which laid the foundation for analyzing sequential data, were commonly utilized for time series forecasting. Exponential smoothing was introduced by [17], [99], and [263] as a method to forecast future values using a weighted average of past data. This method operates by computing an exponential weight decay of historical data, assigning greater weight to more recent data. The Autoregressive Integrated Moving Average (ARIMA) model was formalized through the work of George Box and Gwilym Jenkins in their book “Time Series Analysis: Forecasting and Control” [15]. The ARIMA model predicts future values by leveraging the autocorrelation in data and is composed of three main components. AutoRegressive (AR) uses a linear combination of past values to predict the current value. Moving Average (MA) employs a linear combination of past error terms to predict the current value. Integrated (I) removes non-stationarity (the property where mean and variance change over time) by differencing the data to achieve stationarity. While exponential smoothing models are advantageous for real-time data analysis due to their simplicity, ARIMA models are better suited for capturing complex patterns, making them ideal for long-term forecasting. The SARIMA model extends ARIMA by incorporating seasonal differencing, along with seasonal autoregressive and moving average components, allowing it to effectively model and predict data with regular cyclical patterns. These statistical models are based on specific assumptions and are simple, intuitive, and useful for identifying basic patterns in data.
@@ 3.1.2 Machine Learning Models
3.2 Fundamental Deep Learning Models
@@ 3.2.1 MLPs: The Emergence and Constraints of Early Artificial Neural Networks
Recurrent Neural Networks (RNNs) [100] emerged, opening up new possibilities for processing time series data. RNNs are specialized models designed to process sequential data, such as time series data, and were utilized to overcome the limitations of MLPs. The structure of an RNN consists of an input layer, a hidden layer, and an output layer, and it uses the output from previous time steps(hidden layer output) as the input for the current time step. This structural feature allows RNNs to naturally model the dependencies in data over time. The hidden state preserves past information by storing data from the previous step and passing it on to the next step, thereby maintaining historical context. Additionally, it is parametrically efficient because the same weights are used for calculations across all time steps, regardless of the sequence length.
@@
To address the aforementioned issues with vanilla RNNs, researchers began developing various models. Notably, the Long Short-Term Memory (LSTM) [98] model was developed to address the long-term dependency issues of vanilla RNNs. With the cell state and gate mechanisms (input, output, and forget gates), LSTMs preserve crucial information over extended periods and discard unnecessary information. The Gated Recurrent Unit (GRU) [40] emerged as a simpler alternative to LSTM, offering similar performance while using a more straightforward structure(update and reset gates) to manage information. These two models effectively addressed the vanishing gradient problem and sparked a boom in RNN-based models for time series analysis. Many subsequent RNN-based models used these two as their backbone, making them powerful tools for handling time series data until the advent of Transformers.
@@
The Dilated RNN [25] proposed Dilated Recurrent Skip Connections as a solution to various issues inherent in fundamental vanilla RNNs. This approach enhanced parameter efficiency in long sequences and enabled the network to effectively learn complex dependencies across diverse temporal scales. The DA-RNN [199] attempted to effectively address long-term dependencies, which were challenging for existing methods, by employing a dual-stage attention mechanism. Utilizing an encoder with Input Attention and a decoder with Temporal Attention, it adaptively selected salient information from input features and temporal steps, thereby enhancing the accuracy and performance of time series predictions. Additionally, the MQ-RNN [262] combined the Sequence-to-Sequence neural network framework with Quantile Regression to provide probabilistic forecasts, thereby reflecting uncertainties across various scenarios. This model simultaneously generated multiple forecast horizons at each time step and employed the Forking-Sequences training method to enhance the stability and performance of the RNN.
@@ 3.2.3 CNNs: Extracting key patterns in time series data beyond just images
CNN and RNN Hybrid Models
Building on the popularity of fundamental time series forecasting models like RNNs and the emerging interest in CNN models, hybrid CNN-RNN models began to emerge, combining the strengths of both architectures. Models that combined CNNs and LSTMs were particularly advantageous for simultaneously learning local patterns and long-term dependencies in time series data. The structure involved using CNNs to extract complex features from the time series data and LSTMs to learn temporal dependencies.
DCRNN [134] modeled traffic flow as a diffusion process on a directed graph, capturing both temporal and spatial dependencies. The Diffusion Convolution leveraged bidirectional random walks on the graph to capture spatial dependencies within the traffic network, while the Gated Recurrent Units (GRU) modeled temporal dependencies. The introduction of the Diffusion Convolutional Gated Recurrent Unit (DCGRU) enabled the effective processing of temporal information. TPA-LSTM [223] combined CNN and LSTM with an added attention mechanism, introducing the Temporal Pattern Attention (TPA) methodology to learn important patterns in time series data. Local patterns were extracted using CNN, while the combination of LSTM and the attention mechanism captured significant temporal patterns.
@@ 3.2.4 GNNs: Structurally modeling relationships between variables
Until the development of Transformers, GNNs consistently garnered interest, showing consistent yet slow growth compared to RNNs and CNNs. The development of Graph Convolutional Networks (GCNs) [124] marked another turning point for GNNs. GCNs learn node features through convolution operations on graph structures and are powerful tools for processing graph data. This advancement positioned GCNs as a versatile solution for addressing a wide range of graph-based challenges.
@@
After the era of traditional statistical techniques and machine learning methodologies, fundamental deep learning approaches such as RNNs, CNNs, and GNNs were widely used for time-series analysis and prediction over a long period. However, the advent of Transformers brought a revolutionary change to the AI field. Upon their introduction, Transformers outperformed all existing methods across various domains and continue to dominate as the leading foundation model. In the time-series domain, a similar trend was observed. Beginning with the introduction of the Logsparse Transformer [133], numerous transformer-based models emerged, achieving state-of-the-art performance. As a result, fundamental deep learning methods experienced a period akin to the Ice Age. The remarkable historical TSF models mentioned above are organized chronologically in Fig. 8.
@@ 3.3 The Prominence of Transformer-based Models
Transformer [248], is an innovative model designed to perform complex sequence-to-sequence tasks in natural language processing. This model features an encoder-decoder structure and utilizes a self-attention mechanism to capture the relationships between tokens in the input sequence.
Modeling approaches in self-attention can be broadly categorized into unidirectional and bidirectional paradigms. Unidirectional modeling restricts the self-attention mechanism to consider only the tokens preceding the current token in the sequence. For a sequence $(x_{1},x_{2},\ldots,x_{L})$, the attention for token $x_{t}$ is computed over the subsequence $(x_{1},x_{2},\ldots,x_{t})$ to ensure causality, where $L$ denotes the length of the input sequence. The unidirectional modeling is predominantly employed in tasks requiring sequential generation, such as language generation [18, 46]. In contrast, bidirectional modeling allows the self-attention mechanism to utilize the entire sequence. For token $x_{t}$, attention is computed over $(x_{1},x_{2},\ldots,x_{L})$, thereby leveraging both preceding and succeeding tokens. The bidirectional modeling is employed in tasks requiring holistic context, such as masked token prediction [49, 41].
As a result, Transformers have gained significant attention for replacing fundamental RNN-based models, offering parallel processing capabilities, and effectively addressing long-term dependency issues. Transformers’ success with sequential data naturally led to their extension into time series applications. The attention mechanism plays a significant role in TSF by considering the relationships across all time steps simultaneously, giving more weight to important moments, and effectively learning the trends and patterns in the data.
@@ 3.3.1 Transformer Variants
In recent years, substantial research has focused on transformer-based time series analysis, particularly long-term time series forecasting (LTSF), which has led to the development of various models [294]. The original Transformer model has several limitations when applied to long-term time series forecasting (TSF). These limitations include quadratic time and memory complexity, as well as error accumulation caused by the auto-regressive decoder design. Specifically, the time and memory complexity of self-attention increases quadratically with the length of the input sequence, denoted as $O(L^{2})$, where $L$ represents the length of the input sequence.
This high complexity can become a computational bottleneck, especially for time series forecasting tasks that rely on long-term historical information. To address these issues, Transformer variants primarily focus on reducing time and memory complexity. An attention pattern determines which tokens a token attends to. Fig. 9 compares Full Self-Attention and Sparse Self-Attention patterns to illustrate the efficiency of Sparse Attention [12]. In Fig. 9(a), all tokens attend to each other, sharing global information. In contrast, Fig. 9(b) shows that only a subset of tokens is attended to, reducing computational cost while preserving essential information.
@@
Figure 9: Comparison of Full and Sparse Self-Attention Patterns
LogTrans [133] leverages the efficiency of Log Sparse attention mechanisms to reduce time and memory complexity to $O(LlogL)$, making it well-suited for long sequence forecasting tasks. This model assumes recent data is more relevant than older data, utilizing positional sparsity to focus attention only on relevant data. Reformer [125] enhances efficiency through locality-sensitive hashing and reversible residual layers, reducing memory and computational requirements to $O(LlogL)$ while maintaining performance on long sequences. Locality-sensitive hashing (LSH) approximates attention by clustering similar items and performing attention partially on these clusters. Informer [310] improves computational and memory efficiency to $O(LlogL)$ by utilizing a ProbSparse self-attention mechanism and a generative-style decoder. It effectively handles long sequences and captures dependencies across various time scales. Informer addresses the limitations of Reformer, which performs well only on extremely long sequences, and LogTrans, which uses heuristic positional-based attention. ProbSparse self-attention features dynamic query selection based on query sparsity measurement. Autoformer [269] integrates time series decomposition into the internal blocks of the Transformer and utilizes an auto-correlation mechanism to aggregate similar sub-series. Autoformer improves upon previous methods by addressing point-wise dependency and aggregation issues. It employs a decomposition architecture to handle complex temporal patterns by separating seasonal and trend components. Pyraformer [150] introduces a novel pyramid attention mechanism designed to efficiently model long-term dependencies in time series data. By using a hierarchical structure to progressively reduce sequence length, Pyraformer significantly lowers computational complexity compared to fundamental Transformer models. Fedformer [311] combines Fourier-enhanced decomposition blocks with frequency domain sparse representations, integrating seasonal-trend decomposition from Autoformer and offering a more effective representation of time series data. The sparse representation basis can be Fourier-based or Wavelet-based, encompassing a subset of frequency components, including both high and low frequencies. Additionally, attempts have been made to overcome the limitations of normalized Transformers, which often produce indistinguishable attention patterns and fail to capture significant temporal dependencies. Non-stationary Transformer [151] identifies this issue as excessive normalization and addresses it with De-stationary Attention, which allows for the use of non-normalized data during QKV computations.
@@ 3.3.2 Limitation of Transformer-based Models
Context Window Limited to Current Input
In principle, Transformers can handle inputs of arbitrary length, provided there are sufficient computational resources. However, transformers must process the entire context at once in a single input, causing the memory and computation requirements to grow substantially with longer context length. By contrast, RNNs and State Space Models (SSMs) maintain hidden states over time, allowing them to preserve context beyond the immediate input. We highlight this practical limitation, which often arises when handling very long contexts.
@@ 3.4 Uprising of Non-Transformer-based Models
Contemporary TSF tasks increasingly require the prediction of diverse multivariates and long sequences. Unlike earlier tasks that involved relatively short sequences, the limitations of transformer-based models have become more pronounced when dealing with long sequences. Researchers adhering to the Transformer’s philosophy have begun focusing intensively on two aspects to address these issues: reducing computational complexity and enhancing long-term dependency learning. Conversely, some researchers have reverted to non-transformer methodologies, renewing interest in fundamental deep learning models such as RNNs, CNNs, GNNs, and MLPs. Non-transformer-based models have the potential to overcome the limitations of Transformers, and research has begun to leverage these strengths to move beyond the hegemony of the Transformer backbone.
RNN-based models sequentially process input data, retaining the previous state at each time step to predict the next state. This enables the natural modeling of long-term dependencies and helps overcome the limitations of Transformers by leveraging time-ordering learning. They require relatively less data and memory, making them less restrictive in their application. Additionally, their ability to handle real-time streaming data enhances their versatility across various applications.
@@
This claim gained further traction through subsequent studies, leading to a more diversified interest in backbone models across various architectures. The emergence of this model sparked skepticism regarding the performance of Transformer-based models and served as a catalyst for the exploration of various alternative architectures. The era of Transformer prevalence has come to an end, ushering in a renaissance in time series research, characterized by a reevaluation of fundamental DNNs, the emergence of novel models like Mamba, and new approaches such as the utilization of LMMs.
@@ 4.1 Overcoming Limitations of Transformer
As previously mentioned in 3.4, the performance of simple linear models in LTSF-Linear [294] surpassing fundamental transformer-based models has raised doubts about the effectiveness of Transformers. However, in the field of NLP, Transformers still demonstrate superior performance in handling long-term dependencies in sequential data compared to other models [188]. This observation suggests that while Transformers have great potential, researchers have not fully leveraged their capabilities in time series analysis.
@@ 4.1.1 Patching Technique
Transformers were originally developed for natural language processing (NLP), and applying them to time-series analysis requires appropriate adjustments to fit the domain of time series. Different from the rich semantic information carried by individual word tokens in NLP, individual data points in time series are often similar to their neighboring values, lacking substantial information [178]. Therefore, the point-wise attention mechanism of fundamental models fails to consider the broader context or patterns that may span multiple consecutive time steps and only calculates attention for individual time steps, which makes it difficult to capture the characteristics of time series data. For this reason, considering the surrounding context along with a single time point provides more information in time series data.